OpenStudy (anonymous):

help please :) What is the domain and range of (4x+3)/x^2-1

OpenStudy (anonymous):

the domain is all real numbers but -1 and 1 right??

OpenStudy (usukidoll):

yeah since there is a fraction. we have a restriction so \[(x^2-1) =(x+1)(x-1) \] x=1, -1 so our domain should be all reals except when x =-1 or x = 1. Those restrictions create a blank space in the graph

OpenStudy (anonymous):

what about range?

OpenStudy (usukidoll):

range is in the y-axis... hmm.

OpenStudy (usukidoll):

range is always a pain >_< but here's the graph of the function

OpenStudy (anonymous):

i don't get it :/

OpenStudy (usukidoll):

we have the domain... which x is all reals except when x =1 and x =-1

OpenStudy (usukidoll):

we got the graph of the function ... the range is in the y-axis so it must be all reals except when y is some number ?!

OpenStudy (usukidoll):

@jim_thompson5910 I need to refresh my memory. Range is in the y-axis but how do we find it again when there are restrictions in the function ?

jimthompson5910 (jim_thompson5910):

degree of denominator = 2 degree of numerator = 1 since (degree of denominator) > (degree of numerator), this means y = 0 is a horizontal asymptote. So y = 0 would be kicked out of the range but y = 0 is certainly possible. Just let x = -3/4

OpenStudy (usukidoll):

?!

jimthompson5910 (jim_thompson5910):

plug in x = -3/4 and you'll get y = 0

jimthompson5910 (jim_thompson5910):

so the range is the set of all real numbers

OpenStudy (usukidoll):

how is the range all reals? I don't get that part..

OpenStudy (usukidoll):

oh wait.. unless the line is going through all y -values. ._.

jimthompson5910 (jim_thompson5910):

yeah any y value is possible

OpenStudy (anonymous):

wait i don't get it

OpenStudy (usukidoll):

it is all reals XD . My mind went blank . Why is domain easier to find than range?

OpenStudy (usukidoll):

in the graph the lines is going through all the y-values where as in the x-values, there is a break (like an empty space) at x=1, x = -1

OpenStudy (anonymous):

oh!! i see it, thank you :) can you also help me on finding the relative extrema for 2x^3+5x^2-25x?

OpenStudy (usukidoll):

alright but first we have to graph this

OpenStudy (anonymous):

can we find it without graphing

OpenStudy (usukidoll):

relative extrema.. like maxs and mins.. (gonna be hard though without a graph )

OpenStudy (usukidoll):

there's the first derivative test though for increasing and decreasing.. finding extremas... but I'm not sure if you're at that level yet.

OpenStudy (anonymous):

nope haha im confused. i'm supposed to find it without equation i think

OpenStudy (anonymous):

without the graph*

OpenStudy (usukidoll):

so you're not in a calculus course?

OpenStudy (anonymous):

its precalc!

OpenStudy (usukidoll):

ok... do you know how to take derivatives?

OpenStudy (usukidoll):

weird you need those

OpenStudy (usukidoll):

this site has an example for finding relative extrema http://ltcconline.net/greenl/courses/105/theoremsrelatedrates/extrema.htm

OpenStudy (anonymous):

i figured it out haha thanks :)